# Exploring the Interior Structure and Mode of Tidal Heating in Enceladus

### v3

@@ 2.3 Physical Libration
Owing to Enceladus’s moderate orbital eccentricity and slightly elongated shape, it is subject to periodic torques that force harmonic oscillations in its orientation, called physical librations. This libration is superimposed on the overall synchronous orbit of the moon around Saturn. The amplitude of physical libration depends mainly on the moments of inertia of the body, and whether the outer surface layer, i.e., the ice shell, is coupled to the deep interior (e.g., Thomas et al., 2016), and to a lesser extent to the rigidity of the shell. The estimated libration of Enceladus based on the analysis of Cassini observations is too large to be explained by an ice shell that is mechanically coupled to the core, implying a global ocean (Thomas et al., 2016). Recent work by Park et al. (2024) suggests a slightly lower amplitude of libration, although still higher than expected if the ice shell and the core are coupled. We note that the shell and core might decouple without a global ocean in the presence of a frictionless interface, such as found in soft-bedded glacier systems on Earth (Johnston & Montési, 2017). While the scenario with no global ocean cannot be ruled out by the constraint from libration, here we assume such an ocean is present.
Tides can influence the amplitude of libration by various mechanisms: by modifying the gravitational torque exerted by the planet on the satellite; through gravitational coupling of the libration of the ice shell to the solid interior through the torque between the periodic tidal bulge of one solid layer and the permanent shape of another; and by the slight periodic modification of the polar moment of inertia by the zonal tides.The amplitude of libration of the outer ice shell is given by (Van Hoolst et al., 2013):
@@
where $K_{1}$, $K_{2}$, $K_{3}$, $K_{4}$, $K_{5}$ and $K_{6}$ describe the torques between the planet, the shell, and the core, $C_{c}$ and $C_{s}$ are the polar moments of inertia of the core and the shell, $\gamma$ is the amplitudes of libration of the shell, $\mathcal{M}$ is the mean anomaly, and
@@
Equation (8) represents the natural frequencies of free libration of the body (Van Hoolst et al., 2013). To compute the flattening of each of the layers, as required for the calculation of the torques between layers in equations (78), we use the approach of Tricarico (2014) that provides hydrostatic equilibrium figures for moons in eccentric orbits, along with the global shape from the best-fitting ellipsoid model suggested by Park et al. (2024).
@@ 2.4 Static Gravity Field
Current gravity data for Enceladus is limited to that obtained from three Cassini fly-bys. Therefore, we are limited to gravity field coefficients of spherical harmonic degree $l=2,3$. To further constrain the interior structure, we use the moment of inertia and the $l=3$ zonal potential coefficients ($J_{3}$). The moment of inertia is derived from the gravity coefficients ($J_{2}$ and $C_{22}$) with assumptions to account for the non-hydrostatic component of the gravity field and the compensation mechanism, as discussed in Hemingway & Mittal (2019). Directly measuring the moment of inertia of Enceladus is possible only by measuring the precession of the body. Hemingway et al. (2018) propose to derive the moment of inertia from long-wavelength $l=2$ gravity-topography admittance measurements ($Z_{22}$ and $Z_{20}$). See McKinnon (2015), Hemingway et al. (2018), and Hemingway & Mittal (2019) for details. Here, we only consider the moment of inertia when we explore the potential impact of future observations. To compute the zonal non-hydrostatic degree-3 static gravity coefficient ($J_{3}$), we follow an approach similar to that described in Genova et al. (2024). The gravity-topography admittance is computed as (Turcotte et al., 1981; Hemingway & Matsuyama, 2017):
@@ 3 End-member scenarios for tidal dissipation in Enceladus
The viscosity of consolidated rock is higher than $10^{18}$ Pa.s Hirth & Kohlstedt (e.g., 2003). However, unconsolidated material, or even a mixture of rock and ice, may have a significantly lower effective viscosity. In this context, Choblet et al. (2017) propose a range of lower core viscosities that allows substantial tidal heating in the core of Enceladus. Alternatively, tidal dissipation in the icy crust may explain the tidal heating at Enceladus if the temperature of the ice is sufficiently warm and the effective shell viscosity sufficiently low. The estimated tidal heating from these different scenarios is shown in Figure 1(a). The four narrow regions determined by the red lines indicate the ranges of the estimated tidal heating in Enceladus and the associated effective viscosities of the shell and the core, assuming that the tidal heating is fully balanced by the rate of heat loss at the surface in the range of 15–40 GW (Howett et al., 2011; Hemingway & Mittal, 2019; Nimmo et al., 2023; Park et al., 2024). We show the variation in tidal heating with the viscosity of the shell (core) in Figure 1 panels (b) and (c) for a fixed value of core (shell) viscosity. The dissipation peaks at certain viscosities of both the shell and core, at which the associated Maxwell relaxation time is close to the period of excitations and the tidal response transitions from anelasticity-dominated behavior to viscous-dominated behavior. Therefore, the estimated tidal dissipation rate is compatible with two regions around the peak dissipation (highlighted in gray). For the interior structure assumed here, we find an effective core viscosity O(1013) Pa.s can satisfy the total estimated heat, consistent with that suggested by Choblet et al. (2017).
@@
The two end-member scenarios manifest themselves differently in the amplitudes of $k_{2}$, $h_{2}$, and $l_{2}$ and their respective phase lags (Figure 2). When the effective viscosity of the rocky core is $10^{10}~\rm Pa.s\lesssim\eta_{core}\lesssim~10^{13}$ Pa.s., deformation of the core will significantly affect the amplitude of  $k_{2}$ and its phase lag $\epsilon_{k_{2}}$ (Figure 2, panels (a) and (d)). However, tidal deformation of the core does not substantially affect tidal displacement at the surface, resulting in only a small change in $h_{2}$. i.e., less than 10% for Enceladus. Tidal deformations in the core can have a substantial effect on surface deformations ($h_{2}$) if the viscosity of the shell is very low. This scenario is unlikely due to the low temperatures in most of the ice shell.
@@
The compatibility of the produced tidal heat with different regimes of tidal activity in the shell or in the core (Figures 1 and 2) presents an ambiguity in determining the distribution of tidal dissipation that requires additional constraints to resolve. In the case of Enceladus, and for most of the icy bodies where the ice shell is mechanically decoupled from the solid interior by an ocean, the ambiguity in the location of the tidal heating can be resolved by considering the tidal response at the surface and the associated phase lag. In the case of a mechanically coupled interior or tidally inactive interior layers, there is a strong correlation between the amplitude and phase lag of deformations ($|h_{2}|$, $|l_{2}|$, $\epsilon_{k_{2}}$, $\epsilon_{l_{2}}$) on the surface and those of the observed gravitational signal. However, a lack of such a correlation would imply that tidal activity in deeper interior layers that are decoupled from the surface—such as the core—is causing a perturbation in the body’s gravitational field. If the ice shell is the primary locus of tidal heat generation, then it undergoes considerably larger deformations that incorporates viscoelastic deformations and a significant phase lag, as opposed to the case where the deformations in the shell remain largely elastic. For a given interior structure model of Enceladus, the tidal deformations on the surface in the two cases differ by approximately a factor of two (Figure 2 panels b and c). The phase lags associated with the tidal displacements can differ by tens of degrees between the two cases. Section 6 explores the details of determining the primary location of tidal dissipation.
@@
Ermakov et al. (2021) and Genova et al. (2024) propose suites of geophysical measurements to study the interior of Enceladus. However, the ranges of elastic and viscoelastic parameters that they consider for the rocky core are such that the tidal deformation and dissipation in the core are effectively negligible. Ermakov et al. (2021) assume $\eta_{core}>10^{20}$ Pa.s, $\mu_{core}>10$ GPa, and Genova et al. (2024) assume $\eta_{core}>10^{16}$ Pa.s, $\mu_{core}>50$ GPa. The results of these studies implicitly represent only one of the tidal dissipation scenarios, i.e., tidal dissipation occurring in the shell without an effective contribution from the core. Here, we broaden the parameter space to allow for the possibility of significant tidal dissipation in the core. Note that the commonly assumed viscosity of ice, as suggested by Goldsby & Kohlstedt (2001) ($\eta_{\rm ice}\approx 10^{14}$ Pa.s,) corresponds to ice near its melting point. Due to the strong temperature dependence of viscosity and very low temperature conditions at Enceladus that may persist deep within the shell, a uniform viscosity of $10^{14}$ Pa.s for the shell (Efroimsky, 2018; Ermakov et al., 2021; Genova et al., 2024) is unrealistic.
@@
Figure 1: (a) Tidal heating (in GW) versus the effective viscosity of the ice shell and the core. The narrow regions bounded by the red lines indicate the range of heat loss that covers the estimates by Howett et al. (2011); Hemingway & Mittal (2019); Park et al. (2024). (b) Tidal heating variation with ice shell effective viscosity for a fixed core viscosity of 1020 Pa.s. (c) Tidal heating variation with core effective viscosity for a fixed shell viscosity of 1016 Pa.s. The gray areas highlight the region where tidal heating values are compatible with the estimated steady state heating. The results are presented for $D_{\rm shell}=14$~km, $D_{\rm ocean}=34$ km, $\rho_{\rm ocean}=1000~\rm kg/m^{3}$, $R_{\rm core}=204$ km, $\rho_{\rm shell}=930~\rm~kg/m^{3}$, $\mu_{\rm shell}=3$ GPa, $\kappa_{\rm shell}=10$ GPa, $\rho_{\rm core}=2175~\rm~kg/m^{3}$, $\kappa_{\rm core}=20$ GPa, $\mu_{\rm core}=15$ GPa. We assume an Andrade viscoelastic rheology.
@@
Figure 2: Tidal Love number amplitudes (panels a, b, c), phase lags (d, e, f) for different shell and core viscosities. Same interior structure parameters as in Figure 1 are used.
@@
As described in Section 2.1, various viscoelastic models exist based on theoretical and laboratory results. Because the nominal values of the effective relaxation time of ice could be close to the orbital period (tidal forcing period) of Enceladus, we examine the choice of Maxwell and Andrade viscoelastic models by computing the tidal response components (Love numbers and heat) using both models. Our analysis shows that if the viscosity of the ice shell is sufficiently high due to the low temperature ($\eta_{\rm shell}\gtrsim 10^{18}$ Pa.s), its relaxation time will be significantly longer than the forcing period, implying a predominantly elastic behavior of the shell, and therefore the choice of the viscoelastic model does not significantly affect the response (See Appendix A). However, if the viscosity of the ice shell is such that the relaxation time ($\tau_{M}~=~\eta_{\rm shell}/\mu_{\rm shell}$) is close to the forcing period, that is, $\eta_{\rm shell}~\approx~10^{13-14}$ Pa.s, the Maxwell and Andrade viscoelastic models predict very different surface deformation ($\sim$3 m, see Appendix A). The two viscoelastic models differ more significantly in their predicted imaginary component of the response, and thus in the generated frictional heat. The real part of the response is largely determined by the rigidity of the body. The difference in total tidal heating generated in the ice shell and the core can reach several GW. Although we do not rule out the applicability of the Maxwell viscoelastic model (see e.g., Hussmann et al., 2016), for further calculations in this paper, we adopt the Andrade model because of its ability to capture transitional anelastic mechanisms missing in the Maxwell model. The extent to which Enceladus’s tidal response deviates from a Maxwell body to an Andrade or more complex viscoelastic body is determined by the large-scale rheological state of its interior (see, e.g., Bagheri et al. (2022b); Bierson (2024) for discussions on various viscoelastic models). We investigate the sensitivity of Enceladus’s tidal response to the choice of Andrade parameters in Appendix A.
4 Sensitivity of physical libration to the shell properties
[…]
Figure 3: Variations of (a) Period of natural (free) libration and (b) amplitude of forced libration of Enceladus as a function of shell thickness and density. The orbital period of Enceladus is shown by gray line in (a). The end-member values of amplitude of libration inferred by Thomas et al. (2016); Nadezhdina et al. (2016); Park et al. (2024) (370 $\gamma<~750 m$) are shown by red lines in (b). Other interior structure parameters are the same as in Figure 2. The x and y axes and the libration amplitude color scale are logarithmic.
[…]
@@ 4 Sensitivity of physical libration to the shell properties
Table 2: Prior ranges for interior structure model parameters. We sample all the model parameters on logarithmic scale.
5 Constraining the partitioning of tidal heating between the shell and the core
The total tidal energy dissipated in Enceladus can be computed by measuring the amplitude and phase lag of $k_{2}$, independent of the choice of rheological model (see Equation (5)). We can compute the tidal heating in any specific region of the satellite, such as the core or the shell, by integrating the radial shear dissipation in that part of the body. Here, we use equations (33), (35), and (37) in Tobie et al. (2005) to derive the contribution of the core and the shell to the global tidal heating:
Here, $\hat{k}_{2}^{core}$ is the piecewise gravitational Love number associated with the core. Thus, if we compute the contribution of each of the layers to $|k_{2}|/Q$ or identically to $\Im(k_{2})$ of the body, that is, evaluating at each interface, we can compute the associated tidal heating for the part of the body that is interior to the interface. For a homogeneous core, $\hat{k}_{2}^{core}$ can be computed from the deformation at the top of the core ($\Delta R_{core}$):
$\Im{(\hat{k}_{2}^{core}})=\frac{4\pi G\rho_{core}R_{core}^{4}}{5R_{E}^{3}}\Im{(\Delta R_{core})^{4}},$ (13)
Details of this derivation are presented in Appendix C. The tidal heating in the ice shell is then calculated by subtracting the contribution of the core from the total energy dissipated:
$\dot{E}_{shell}=\dot{E}-\dot{E}_{core}.$ (14)
Tidal dissipation in the ocean is assumed negligible (Hay & Matsuyama, 2017; Matsuyama et al., 2018)
Equations (12) and (14) provide a method to calculate the partitioning of the global tidal heating for a given interior structure model. These equations require a priori assumptions or independent constraints on the rheological and structural properties to compute the shear dissipation in the core. Therefore, here we derive the partitioning of tidal heating on the basis of the observable quantities, i.e., the tidal Love numbers. We extend the formulation provided by Beuthe (2019) which uses membrane mechanics to obtain the contribution of each layer of the body to the imaginary part of the tidal Love number. The shell’s contribution can be written as (Beuthe, 2019):
where, following the terminology used in Beuthe (2019), $\xi_{2}=3\rho_{ocean}/5\rho_{E}$, $\rho_{E}$ is the density of the whole body, $k_{2}$ and $h_{2}$ are the tidal Love numbers of the whole body, and $k_{2}^{0}$ is the fluid-crust gravitational Love number (assuming the crust is fluid), and $\Lambda_{T}$ is the shell spring constant introduced by Beuthe (2019). In Appendix C, we show that the tidal dissipation occurring in the shell can be expressed in terms of the tidal Love numbers and density of the layers:
Here, $\delta\rho$ is the density difference between the ocean and the crust. $\epsilon$ is the ratio of the thickness of the crust to Enceladus’s radius. Equation (16) expresses the tidal heating in the shell using measurements of the real and imaginary parts of the Love numbers ($k_{2}$, $h_{2}$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$), and information on the density structure. The densities of the layers of the body can be constrained using the static gravity field and libration, and are easier to constrain compared to the rheology of the layers required in Equation (15). Based on Equation (16), the uncertainty in the magnitude of tidal heating occurring in the shell can be computed from the uncertainties on the density structure and the complex Love numbers. The tidal heating in the core can be computed by subtracting the tidal dissipation in the shell from the total heat. This derivation is general and can be used for all other icy satellites for which the necessary geodetic observations are or will become available.
6 Road-map to Constrain the interior properties

### v5

@@ 2.4 Physical Libration
Owing to Enceladus’s moderate orbital eccentricity and slightly elongated shape, it is subject to periodic torques that force harmonic oscillations in its orientation, called physical librations. This libration is superimposed on the overall synchronous orbit of the moon around Saturn. The amplitude of physical libration depends mainly on the moments of inertia of the body, and whether the outer surface layer, i.e., the ice shell, is coupled to the deep interior (e.g., Thomas et al., 2016), and to a lesser extent to the rigidity of the shell. The estimated libration of Enceladus based on the analysis of Cassini observations is too large to be explained by an ice shell that is mechanically coupled to the core, implying a global ocean (Thomas et al., 2016). Recent work by Park et al. (2024) suggests a slightly lower amplitude of libration, although still higher than expected if the ice shell and the core are coupled. We note that presence of a frictionless interface, such as found in soft-bedded glacier systems on Earth (Johnston & Montési, 2017) might also decouple the shell from the core. While the scenario with no global ocean cannot be ruled out by the existing constraints, here we assume the presence of a global subsurface ocean.
Tides can influence the amplitude of libration by various mechanisms: modifying the gravitational torque exerted by the planet on the satellite; gravitational coupling of the libration of the ice shell to the solid interior through the torque between the periodic tidal bulge of one solid layer and the permanent shape of another; and the slight periodic modification of the polar moment of inertia by the zonal tides.The amplitude of libration of the outer ice shell is given by (Van Hoolst et al., 2013):
@@
where $\gamma$ is the amplitudes of libration of the shell, $K_{1}$, $K_{2}$, $K_{3}$, $K_{4}$, $K_{5}$ and $K_{6}$ describe the torques between the planet, the shell, and the core, $C_{c}$ and $C_{s}$ are the polar moments of inertia of the core and the shell, $\mathcal{M}$ is the mean anomaly, and
@@
$\sigma_{1,2}$ represents the natural frequencies of free libration of the body (Van Hoolst et al., 2013). To compute the flattening of each of the layers, as required for the calculation of the torques between layers in equations (12, 13), we use the approach of Tricarico (2014) that provides hydrostatic equilibrium figures for moons in eccentric orbits, along with the global shape from the best-fitting ellipsoid model suggested by Park et al. (2024).
@@ 2.5 Static Gravity Field
Current gravity data for Enceladus is limited to that obtained from three Cassini fly-bys. Therefore, we are limited to gravity field coefficients of spherical harmonic degrees $l=2,3$. To further constrain the interior structure, we use the moment of inertia and the $l=3$ zonal potential coefficients ($J_{3}$). Direct measurement of the moment of inertia of Enceladus is only possible by measuring the body’s precession. Alternatively, Hemingway et al. (2018) propose to derive the moment of inertia from long-wavelength $l=2$ gravity-topography admittance measurements ($Z_{22}$ and $Z_{20}$). See McKinnon (2015), Hemingway et al. (2018), and Hemingway & Mittal (2019) for details. This approach requires underlying assumptions about the non-hydrostatic component of the gravity field and the compensation mechanism, as discussed in Hemingway & Mittal (2019). Here, we only consider the moment of inertia when we explore the potential impact of future observations. To compute the zonal non-hydrostatic degree-3 static gravity coefficient ($J_{3}$), we follow an approach similar to that described in Genova et al. (2024). The gravity-topography admittance is computed as (Turcotte et al., 1981; Hemingway & Matsuyama, 2017):
@@
3 Sensitivity of Physical libration to shell properties
[…]
Figure 1: Variations of (a) Period of natural (free) libration and (b) amplitude of forced libration of Enceladus as a function of shell thickness and density. The orbital period of Enceladus is shown by gray line in (a). The end-member values of amplitude of libration inferred by Thomas et al. (2016); Nadezhdina et al. (2016); Park et al. (2024) (370 $\gamma<750 m$) are shown by red lines in (b). The results are presented for $\rho_{\rm ocean}=1000~\rm kg/m^{3}$, $R_{\rm core}=197$ km, $\rho_{\rm shell}=900~\rm~kg/m^{3}$, $\mu_{\rm shell}=3$ GPa, $\kappa_{\rm shell}=10$ GPa, $\eta_{shell}=6\times 10^{13}$ Pa.s, $\rho_{\rm core}=2320~\rm~kg/m^{3}$, $\mu_{\rm core}=40$ GPa, $\kappa_{\rm core}=100$ GPa, $\eta_{core}=10^{19}$ Pa.s. The horizontal and vertical axes and the libration amplitude color scale are logarithmic.
[…]
@@ 4 End-member scenarios for tidal dissipation in Enceladus
In this section, we discuss the two main categories of tidal dissipation and the ranges of the effective viscosities of the shell and core that are consistent with the inferred heating rate. The viscosity of consolidated rock is higher than $10^{18}$ Pa.s (e.g., Hirth & Kohlstedt, 2003). However, unconsolidated material, or even a mixture of rock and ice, may have a significantly lower effective viscosity. Choblet et al. (2017) propose a range of low core viscosities that allows for substantial tidal heating in the core of Enceladus. Alternatively, tidal dissipation in the icy crust may explain the tidal heating the effective shell viscosity is sufficiently low, i.e., if the temperature of the ice is sufficiently high. The estimated tidal heating for broad ranges of shell and core viscosities is shown in Figure 2(a). We assume that the total heat loss is fully balanced by the generated tidal heat computed from equation (5). In other words, we assume that Enceladus is currently in a thermal steady state. We assume a tidal heating rate in the range of 15–40 GW that covers the estimates by Howett et al. (2011); Hemingway & Mittal (2019); Nimmo et al. (2023); Park et al. (2024). Each of the four narrow regions determined by the red lines in Figure 2(a) indicate the ranges of the estimated tidal heating in Enceladus and the associated effective viscosities of the shell and the core. We show the variation in tidal heating with the viscosity of the shell (core) in Figure 2 panels (b) and (c) for a fixed value of core (shell) viscosity. The dissipation peaks at certain viscosities of both the shell and core, at which the associated Maxwell relaxation time is close to the period of excitations and the tidal response transitions from anelasticity-dominated behavior to viscous-dominated behavior. Therefore, the estimated tidal dissipation rate is compatible with two regions around the peak dissipation (highlighted in gray). For the interior structure assumed here, we find that an effective core viscosity O(1013) Pa.s can satisfy the total estimated heat, consistent with Choblet et al. (2017).
@@
The two end-member scenarios manifest themselves differently in the amplitudes of $k_{2}$, $h_{2}$, and $l_{2}$ and their phase lags (Figure 3). When the effective viscosity of the rocky core is $10^{10}~\rm Pa.s\lesssim\eta_{core}\lesssim 10^{13}~\rm Pa.s.$, deformation of the core will significantly affect the amplitude of  $k_{2}$ and its phase lag $\epsilon_{k_{2}}$ (Figure 3, panels (a) and (d)). However, tidal deformation of the core does not substantially affect tidal displacement at the surface, resulting in only a small change in $h_{2}$. i.e., less than 10%. Tidal deformations in the core can have affect the surface deformations ($h_{2}$, $l_{2}$) if the viscosity of the shell is very low. This scenario is unlikely given the low temperatures prevailing through most of the ice shell.
@@
The compatibility of the produced tidal heat with different regimes of tidal activity in the shell or in the core (Figures 2 and 3) presents complexities in determining the distribution of tidal dissipation that requires additional constraints to resolve. In the case of Enceladus, and for most of the icy bodies where the ice shell is mechanically decoupled from the solid interior by an ocean, the ambiguity in the location of the tidal heating can be resolved by considering the tidal response at the surface and the associated phase lag. In the case of a mechanically coupled interior or tidally inactive interior layers, there is a strong correlation between the amplitude and phase lag of surface deformations (represented by $|h_{2}|$, $|l_{2}|$, $\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$) and those of the observed gravitational signal ($k_{2}$ and $\epsilon_{k_{2}}$). A lack of such a correlation would imply that tidal activity in deeper interior layers which are decoupled from the surface—such as the core—is causing a perturbation in the body’s gravitational field. If the ice shell is the primary locus of tidal heat generation, then it undergoes larger deformations and has a significant phase lag, a response that incorporate viscoelastic deformation, as opposed to the case where the deformations in the shell remain purely elastic. For a given interior structure model of Enceladus, the tidal deformations on the surface in the two cases differ by approximately a factor of two (Figure 3 panels b and c). The phase lags associated with the tidal displacements can differ by tens of degrees between the two cases. Section 5 explores the details of determining the primary location of tidal dissipation.
@@
Ermakov et al. (2021) and Genova et al. (2024) propose suites of geophysical measurements to study the interior of Enceladus. However, Ermakov et al. (2021) assume $\eta_{core}>10^{20}$ Pa.s, $\mu_{core}>10$ GPa, and Genova et al. (2024) assume $\eta_{core}>10^{16}$ Pa.s, $\mu_{core}>50$ GPa, the ranges of elastic and viscoelastic parameters that imply that the core is effectively negligible. Therefore, the results of these studies represent only one of the tidal dissipation scenarios, i.e., tidal dissipation occurring in the shell without any effective contribution from the core. Here, we broaden the parameter space to allow for the possibility of significant tidal dissipation in the core. Note that the commonly assumed viscosity of ice, as suggested by Goldsby & Kohlstedt (2001) ($\eta_{\rm ice}\approx 10^{14}$ Pa.s,) corresponds to ice near its melting point. Due to the strong temperature dependence of viscosity and very low temperature conditions at Enceladus that may persist deep within the shell, a uniform viscosity of $10^{14}$ Pa.s for the shell (Efroimsky, 2018; Ermakov et al., 2021; Genova et al., 2024) is unrealistic.
@@
Figure 2: (a) Total tidal heating (in GW) versus the effective viscosity of the ice shell and the core. The narrow regions bounded by the red lines indicate the range of heat loss that covers the estimates by Howett et al. (2011); Hemingway & Mittal (2019); Park et al. (2024). (b) Tidal heating variation with ice shell effective viscosity for a fixed $\eta_{core}=10^{20}$ Pa.s. (c) Tidal heating variation with core effective viscosity for a fixed $\eta_{shell}=10^{16}$ Pa.s. The gray areas highlight the region where tidal heating values are compatible with the estimated steady state heating. The results are presented for $D_{\rm shell}=14$~km, $D_{\rm ocean}=34$ km, $\rho_{\rm ocean}=1000~\rm kg/m^{3}$, $R_{\rm core}=204$ km, $\rho_{\rm shell}=930~\rm~kg/m^{3}$, $\mu_{\rm shell}=3$ GPa, $\kappa_{\rm shell}=10$ GPa, $\rho_{\rm core}=2175~\rm~kg/m^{3}$, $\kappa_{\rm core}=20$ GPa, $\mu_{\rm core}=15$ GPa. We assume an Andrade viscoelastic rheology.
@@
Figure 3: Tidal Love number amplitudes (a, b, c), and phase lags (d, e, f) for different shell and core viscosities. Same interior structure parameters as in Figure 2 are used.
@@
As described in Section 2.1, various viscoelastic models exist based on theoretical and laboratory results. Because the nominal values of the effective relaxation time of ice could be close to the orbital period (tidal forcing period) of Enceladus, we examine the choice of Maxwell and Andrade viscoelastic models by computing the tidal response components (Love numbers and heat) using both models. Our analysis shows that if the viscosity of the ice shell is sufficiently high due to the low temperature ($\eta_{\rm shell}\gtrsim 10^{18}$ Pa.s), its relaxation time will be significantly longer than the forcing period, implying a predominantly elastic behavior of the shell, and therefore the choice of the viscoelastic model does not significantly affect the response (See Appendix A). However, if the viscosity of the ice shell is such that the relaxation time ($\tau_{M}~=~\eta_{\rm shell}/\mu_{\rm shell}$) is close to the forcing period, that is, $\eta_{\rm shell}~\approx~10^{13-14}$ Pa.s, Maxwell and Andrade viscoelastic models predict considerably different surface deformation ($\sim$3 m, see Appendix A). The two viscoelastic models differ more significantly in their predicted imaginary component of the response, and thus in the generated frictional heat. The real part of the response is largely determined by the rigidity of the body. The difference in total tidal heating generated in the ice shell and the core can reach several GW. Although we do not rule out the applicability of the Maxwell viscoelastic model (see e.g., Hussmann et al., 2016), for further calculations in this paper, we adopt the Andrade model because of its ability to capture transitional anelastic mechanisms missing in the Maxwell model. The extent to which Enceladus’s tidal response deviates from a Maxwell body to an Andrade or more complex viscoelastic body is determined by the large-scale rheological state of its interior (see, e.g., Bagheri et al. (2022b); Bierson (2024) for discussions on viscoelastic models). We investigate the sensitivity of Enceladus’s tidal response to the choice of Andrade parameters in Appendix A.
@@
Table 2: Prior ranges for interior model parameters. We sample all the model parameters on logarithmic scale.
5 Road-map to Constrain the interior structure and tidal heating mode